OpenStudy (3psilon):

Why is e^ lnx^3 equal to just x^3?

OpenStudy (lgbasallote):

properties of logarithms... \[\huge e^{\ln a} \implies e^{\log_e a} \implies a\] do you remember this property?

OpenStudy (tkhunny):

e^(ln(x)) = x ln(e^x) = x They are inverse functions.

OpenStudy (anonymous):

e is a inverse Function of ln so they cancles off....like \[\sin(\sin ^{-1} 7) = 7\]

OpenStudy (lgbasallote):

you can;t take the arcsin of 7....

zepdrix (zepdrix):

lol i suppose that's true :) Example gone bad... i hate when that happens XD

OpenStudy (anonymous):

oh...Sorry.....Max value of sin is 1.....) hm...hm